OpenStudy (alicemaombi):

Use the zero product property to find the solutions to the equation x2 – 9 = 16.

OpenStudy (gniiij):

hi

OpenStudy (gniiij):

so what you want to do is isolate the x^2

OpenStudy (gniiij):

to do that you want to add 9 to both sides

OpenStudy (gniiij):

you should get x^2=25

OpenStudy (gniiij):

now square root both sides

OpenStudy (gniiij):

and you get...\[\pm5\]

OpenStudy (peachpi):

@GNiiij answer is correct. To display the zero product property you have to make one side 0, then factor. x² - 25 = 0 (x - 5)(x + 5) = 0 x = ± 5